OpenStudy (anonymous):

(6 + 3sqrt2)(5 - sqrt2)

Parth (parthkohli):

Do you know how to multiply two polynomials?

OpenStudy (anonymous):

\[\LARGE (6+ \sqrt[3]{2}) (5-\sqrt{2})\]

OpenStudy (anonymous):

I do not.

Parth (parthkohli):

Hmm, do you know anything about the distributive property?

OpenStudy (anonymous):

A little.

Parth (parthkohli):

OK, now assume that \(5 - \sqrt 2 = y\). You have now,\[y(6 + \sqrt[3]{2})\]Getting it?

OpenStudy (anonymous):

Yes

Parth (parthkohli):

Can you use the distributive property and distribute \(y\)?

OpenStudy (anonymous):

WHat?